OpenStudy (meowlover17):

Given the functions f(x) = 6x + 11 and g(x) = x + 6, which of the following functions represents f[g(x)] correctly?

OpenStudy (meowlover17):

A. f[g(x)] = 6x + 47 B.f[g(x)] = 6x + 17 C.f[g(x)] = 6x2 + 47 D. f[g(x)] = 6x2 + 17

OpenStudy (meowlover17):

Please help

OpenStudy (meowlover17):

@imqwerty

OpenStudy (anonymous):

Hi MeowLover17.

OpenStudy (meowlover17):

Hello do you know how to solve this?

OpenStudy (meowlover17):

@imqwerty Do you understnad it?

OpenStudy (anonymous):

I'm new here so please excuse me if I do something wrong. What the question is asking you to do is find the value of f(x) for any given g(x). What you need to do in order to solve this is merely plus g(x) in for the x value(s) in f(x). In this example, you can see that when we plug g(x) into f(x) we get 6(x+6) + 11. We can simplify this to get f(g(x) = 6x + 36 + 11 or 6x + 47

OpenStudy (meowlover17):

So the answer would be A?

imqwerty (imqwerty):

f(g(x)) = 6(x+6)+11 = 6x+36+11 = 6x+47

OpenStudy (meowlover17):

Thanks boys :D

OpenStudy (meowlover17):

I have another question @imqwerty Could you help me with that one as well

imqwerty (imqwerty):

ok

OpenStudy (meowlover17):

I'll open a new poll